Hydra Report - May 2002: System work & Eng Results

Updated on June 15, 2021, 7:49 am

Detector, dewar and Arcon work, Ramon Galvez H., June 2002

CCD Dewar

  1. Leak tested.
     
  2. Determine gasification rate = 1.8 SCFH
     
  3. Determine As Is Thermal Radiation Load = 3.8 Watts
     
  4. New vacuum valve installed.

 

CCD & Arcon

  1. Dark Current = 1.43 e/pix/hr ; Conditions = Bin 2x2, Gain 1, all sorrounding rooms and dome lights ON.
     
  2. Dark Current = 0.88 e/pix/hr ; Conditions = Bin 2x2, Gain 1, all lights OFF.
     
  3. New ADC bd prepared + tested ( s/n4 new batch). Got no more missing codes around 15k ADU.
     
  4. Video original and spare bd. Noise comparison : both at 3e.
     
  5. Not a single crash. Neither during exposure tests nor while downloading.

 

Dewar Cryogenics

  1. A new LN2 refill and solidification device was designed. I had the good advices from Brooke on this. Tested in the lab and, along with Daniel, on the telescope.
     
  2. Cryo Status : 24hrs hold time, one daily refill. Cryo system ‘tuned’; as a result the refill is done without necessity of vacuum valve control.